OpenStudy (anonymous):

f(x)=2abs(x-40) g(x)=(1/2)x+40 when x is less than 40 and 60 when x is equal or greater than40 Find the derivative of f(g(x)) at x=70 @Calculus1

OpenStudy (anonymous):

Can i explain you the question?

OpenStudy (anonymous):

Are they on?

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Hw do u delete

hero (hero):

Click your name

OpenStudy (anonymous):

i think i understand the questiona

OpenStudy (anonymous):

Wait nut I gotta explain it

OpenStudy (anonymous):

you have to find the composite function (it will be piecewise) then take the derivative then evaluate

OpenStudy (anonymous):

Really g(x) was a graph and I guessed the equation

OpenStudy (anonymous):

It was an equation of a line (1/2)x+40 until x = 40 and then at x =40 there was a corner and it became a straight line so I guessed that this was the equation

OpenStudy (anonymous):

\[f(x) = 2|x-40| = \left\{\begin{array}{rcc} 80-2x& \text{if} & x < 40\\ 2x -80 & \text{if} & x \geq 40 \end{array} \right. \]

OpenStudy (anonymous):

What I am totally confused

OpenStudy (anonymous):

I need f(g(x))

OpenStudy (anonymous):

i am getting to that. it is complicated by the fact that you have two piecewise functions that you have to find the composition of. that is the hard part

OpenStudy (anonymous):

Well f(x) is peicewise?

OpenStudy (anonymous):

absolute value is called peice wise?

OpenStudy (anonymous):

yes it is defined piecewise. for example \[|x| = \left\{\begin{array}{rcc} -x & \text{if} & x <0 \\ x& \text{if} & x \geq 0 \end{array} \right.\]

OpenStudy (anonymous):

Oh cool way of looking at it

OpenStudy (anonymous):

that is why taking the derivative is a pain. you have to know what interval you are talking about. the function is piecewise and so is its derivativfe

OpenStudy (anonymous):

so that explains why i broke up the original function into two parts. now we can find the composition

OpenStudy (anonymous):

Well I don't treat an absolute value as a peicewise I think of it like this \[abs(x)=\sqrt{x ^{2}}\]

OpenStudy (anonymous):

but now we are almost home free, because you want the derivative at x = 70, and you said \[g(70)=60\] right?

OpenStudy (anonymous):

If the equation has a corner then that means it is peicewise?

OpenStudy (anonymous):

Cuz they gave me the graph and I kind of guessed that it was a peicewise function

OpenStudy (anonymous):

well usually if you have a functions whose picture is not a smooth curve, you have to define it two different ways at least. if you want an equation that it. if you have something that looks like 

OpenStudy (anonymous):

you will not be able to write it as one equation

OpenStudy (anonymous):

ok I will draw the function

OpenStudy (anonymous):

ok but it doesn't matter that much because the derivative of your function will be one of two constants. it will either be 2 or -2 depending on the values of x

OpenStudy (anonymous):